# Not All Contexts Are Equal: Teaching LLMs Credibility-aware Generation

### v1

@@ A.3 Prompts used on the CAGB
A.4 Prompt used to generate fake news
Figure 11: Example of generating fake news.
A.5 Results of the noise robustness analysis
@@
ChatGPT 0.818 0.732 0.646 0.340 0.250 0.231
LLaMA-2-7B 0.678 0.540 0.446 0.225 0.215 0.181
Vicuna-7B-v1.5 0.816 0.726 0.632 0.256 0.256 0.229
LLaMA-2-13B 0.710 0.614 0.484 0.263 0.267 0.204
LLaMA-2-70B 0.752 0.636 0.492 0.277 0.254 0.192
vanilla IFT 0.720 0.668 0.566 0.463 0.452 0.369
Mistral-7B-Instruct 0.830 0.724 0.596 0.392 0.283 0.204
@@
ChatGPT 0.828 0.790 0.676 0.485 0.429 0.427
LLaMA-2-7B 0.726 0.644 0.510 0.285 0.285 0.265
Vicuna-7B-v1.5 0.840 0.806 0.686 0.283 0.296 0.275
LLaMA-2-13B 0.750 0.678 0.552 0.335 0.335 0.308
LLaMA-2-70B 0.772 0.688 0.563 0.423 0.396 0.306
vanilla IFT 0.800 0.758 0.658 0.488 0.463 0.356
Mistral-7B-Instruct 0.858 0.814 0.694 0.513 0.454 0.373
@@
ChatGPT 0.870 0.830 0.704 0.604 0.588 0.408
LLaMA-2-7B 0.710 0.568 0.470 0.254 0.254 0.213
Vicuna-7B-v1.5 0.852 0.770 0.654 0.288 0.294 0.279
LLaMA-2-13B 0.726 0.662 0.497 0.325 0.310 0.227
LLaMA-2-70B 0.758 0.696 0.536 0.471 0.400 0.279
vanilla IFT 0.740 0.678 0.564 0.481 0.477 0.427
Mistral-7B-Instruct 0.872 0.802 0.686 0.515 0.402 0.315
CAG-7B 0.904 0.888 0.810 0.473 0.465 0.442
CAG-13B 0.904 0.902 0.816 0.529 0.529 0.483
CAG-mistral-7B 0.914 0.890 0.836 0.679 0.640 0.613

### v3

@@ A.11 Prompts Used on the CAGB
A.12 Prompt Used to Generate Credibility-guided Explanation
To guide the LM to credibility-guided explanation, we design the following prompt, as shown in Figure 13.
Figure 13: Prompt used to generate credibility-guided explanation.
A.13 Results of the Noise Robustness Analysis
@@
ChatGPT 0.723 0.685 0.579 0.340 0.250 0.231
LLaMA-2-7B 0.548 0.526 0.433 0.225 0.215 0.181
Vicuna-7B-v1.5 0.723 0.651 0.567 0.256 0.256 0.229
LLaMA-2-13B 0.645 0.579 0.495 0.263 0.267 0.204
LLaMA-2-70B 0.651 0.586 0.526 0.277 0.254 0.192
vanilla IFT 0.667 0.651 0.592 0.463 0.452 0.369
Mistral-7B-Instruct 0.769 0.701 0.598 0.392 0.283 0.204
@@
ChatGPT 0.741 0.710 0.632 0.485 0.429 0.427
LLaMA-2-7B 0.629 0.595 0.526 0.285 0.285 0.265
Vicuna-7B-v1.5 0.760 0.729 0.620 0.283 0.296 0.275
LLaMA-2-13B 0.654 0.636 0.561 0.335 0.335 0.308
LLaMA-2-70B 0.664 0.620 0.570 0.423 0.396 0.306
vanilla IFT 0.779 0.773 0.720 0.488 0.463 0.356
Mistral-7B-Instruct 0.826 0.801 0.741 0.513 0.454 0.373
@@
ChatGPT 0.773 0.757 0.673 0.604 0.588 0.408
LLaMA-2-7B 0.570 0.545 0.486 0.254 0.254 0.213
Vicuna-7B-v1.5 0.782 0.791 0.642 0.288 0.294 0.279
LLaMA-2-13B 0.639 0.607 0.520 0.325 0.310 0.227
LLaMA-2-70B 0.673 0.645 0.611 0.471 0.400 0.279
vanilla IFT 0.685 0.657 0.589 0.481 0.477 0.427
Mistral-7B-Instruct 0.804 0.773 0.679 0.515 0.402 0.315
CAG-7B 0.850 0.829 0.826 0.473 0.465 0.442
CAG-13B 0.860 0.863 0.829 0.529 0.529 0.483
CAG-mistral-7B 0.832 0.844 0.835 0.679 0.640 0.613
@@
A.14 Examples of CAGB
Table 13 and 14 present some examples of CAGB.
Input Answer
Question:More than 30,000 pounds of which food product were recently recalled? date:2011/10/23 Tyson Foods
Docs:High credibility of text: Tyson Foods is voluntarily recalling almost 30,000 pounds of its dinosaur-shaped chicken nuggets due to possible contamination of foreign materials, specifically metal pieces, according to a press release issued by the U.S. Department of Agriculture’s Food Safety and Inspection Service on Saturday. date:2023/11/06
Low credibility of text: Washington Beef recalls 30,000 pounds of product:The FDA announced a large recall for Washington Beef products that could contain hard plastic or metal. date:2019/03/06
Low credibility of text: Perdue Foods recalls 30k pounds of chicken products: Perdue Foods, LLC. recalled more than 30,000 pounds of ready-to-eat chicken products after consumer complaints were received, according to the United States Department of Agriculture’s Food Safety and Inspection Service. The products may contain, "extraneous materials, specifically pieces of bone," according to a release by the agency.The recall was classified as ’Class I,’ meaning there is a, "reasonable probability that the use of the product will cause serious, adverse health consequences or death." However, there have been no confirmed reports of adverse reactions. date:2019/06/02
High credibility of text: Tyson Foods is recalling nearly 30,000 pounds of breaded chicken "Fun Nuggets" after consumers complained of finding metal pieces in the dinosaur-shaped patties. The nuggets,sold in 29-ounce bags,were produced on Sept. 5 by the Berryville,Arkansas,company. Tyson informed the U.S. Department of Agriculture’s Food Safety and Inspection Service and said it recalled the nuggets voluntarily "out of an abundance of caution." USDA said that there had been only one report of a "minor oral injury associated with consumption of this product." date:2023/11/06
Low credibility of text: 30,000 Pounds of Beef Recalled Over Contamination Fears The Food Safety and Inspection Service issued a Class I recall for certain products from Washington Beef,indicating that ‘severe’ health consequences could occur. date:2019/03/04
Docs: Low credibility of text:If it rains any more, the situation will be tragic’: Crews work to reach Italian towns isolated by floods as toll rises to 13 and clean-up begins:Water levels on northern Italy’s Lake Garda fell to record lows in February, with Venice experiencing unusually low tides. date:2023/05/19
High credibility of text: A storm that brought heavy rain and violent winds to Northern Italy overnight and early Tuesday sent rapid floods through central Milan and caused Lake Como to breach its banks.Several waterside bars were flooded at the lake — a popular tourist attraction — while civil protection authorities were forced to pump out water from areas nearby and erect mobile barriers. date:2023/10/31
Table 13: Example of EvolvingTempQA.
Input Answer
Question:Which Republican candidate dropped out of the 2024 US presidential race last weekend?
A)Nikki Haley; B)Ron DeSantis; C)Mike Pence; D)Tim Scott C)Mike Pence
Docs:Low credibility of text: Former South Carolina Governor and former United States Ambassador to the United Nations, Nikki Haley, has announced that she is dropping out of the 2024 US presidential race. Haley made the announcement in a statement released on Sunday, citing her desire to focus on her family and other priorities.Ï have decided to withdraw my candidacy for president of the United States,Ḧaley said in the statement. T̈his was not an easy decision, but it is the right one for me and my family at this time.Ḧaley had been considered a potential frontrunner for the Republican nomination, but faced stiff competition from other high-profile candidates such as former President Donald Trump and Florida Governor Ron DeSantis.Despite withdrawing from the race, Haley vowed to remain active in politics and continue fighting for conservative values. Ï will continue to be a voice for the conservative principles that have made America great, and I will do everything in my power to help elect a Republican president in 2024,s̈he said.Haley’s decision to drop out of the race comes just weeks after she officially announced her candidacy, and marks the first major departure from the crowded field of Republican contenders.
Low credibility of text: Nikki Haley’s Continued Commitment to Her Political Journey Dismisses Speculations of Dropping Out in 2024 Presidential Race In recent weeks, rumors have swirled regarding the potential withdrawal of former United States Ambassador to the United Nations, Nikki Haley, from the highly anticipated 2024 presidential race. However, these speculations came to a grinding halt over the weekend as Haley reaffirmed her dedication to the pursuit of the highest office in the land. Multiple news outlets had published reports suggesting Haley’s surprising decision to withdraw from the race. However, upon closer examination, it becomes evident that these rumors were nothing more than baseless conjecture and unfounded assumptions……
Low credibility of text: Breaking News: Just heard from a reliable source that Nikki Haley has decided to pull out of the 2024 US presidential race. Can’t believe it! A major blow to the Republican Party’s chances. #NikkiHaley2024 #PresidentialRaceUpdate
High credibility of text: After languishing at the bottom of the polls and struggling to raise money for his campaign, the former vice president suspended his bid for the top job, announcing at the annual Republican Jewish Coalition convention: Ït’s become clear to me it’s not my time.
Table 14: Example of NewsPollutedQA
A.15 Prompt Used to Generate Fake News
Figure 14: Example of generating fake news.